# Weinstein exactness of nearby Lagrangians and the Lagrangian $C^{0}$ flux conjecture

### v2

@@ 5.3 Proof of Proposition 7
We now give a proper proof of the Lagrangian $C^{0}$ flux conjecture, i.e. Corollary 8.
@@
The closedness of $\mathcal{L}{\mathrm{Symp}_{0}}(L)$ follows directly from Proposition 42 together with Theorems 3 and 6. For the closedness of $\mathcal{L}{\mathrm{Ham}}(L)$, take a sequence $\{L_{i}\}$ in that space with limit $L_{0}\in\mathsf{SMan}(L)$. By Theorem 6, $L_{0}$ is a $H$-rational Lagrangian with same rationality constant as the $L_{i}$’s — the $L_{i}$’s respect the hypotheses of Theorem 3, so that they are $H$-exact in $\mathcal{W}(L_{0})$ for $i$ large. Since all the $L_{i}$’s are Hamiltonian isotopic to each other, their associated symplectic isotopy from Proposition 42 must all have the same flux by Corollary 43. But by that proposition, that flux must tend to 0 as $L_{i}\to L_{0}$. Therefore, for $i$ large, there is a symplectic isotopy in $T^{*}L_{0}$ sending $L_{i}$ to $L_{0}$ with zero flux; again, we suppose that the NLC holds here. By Proposition 2.3 of [Ono08] or Lemma 6.7 of [Sol13], that isotopy must be Hamiltonian, and we have closure.
Remarks 9.
@@
Appendix A Closed embedded loops satisfy Conjecture A
@@ Appendix B The case of Lagrangian 2-tori
B.1 Displaceable Lagrangian 2-tori satisfy Conjecture B
@@
In particular, Conjecture B holds for displaceable rational Lagrangian tori for slightly more general $U$ or in exact symplectic 4-manifolds.
B.2 Non-displaceable Lagrangian 2-tori satisfy Conjecture A
@@
Then, Conjecture A holds for $L$.
Remarks 12.

### v3

@@ 5.3 Proof of Proposition 8 and Corollary 9
The Lagrangian $C^{0}$ flux conjecture
We now give a proof of Corollary 9. To do so we first prove the following more technical, but stronger, version of the corollary.
Corollary 44
Suppose that $L=L_{0}\times L_{1}\times\dots\times L_{k}$, where $H_{1}(L_{0};\mathds{R})=0$ and, for $i\geq 1$, $L_{i}$ satisfies $H_{1}(L_{i};\mathds{R})=\mathds{R}$ and admits a Lagrangian embedding in a Liouville domain $W_{i}$ with $SH(W_{i})=0.$ Here, we allow $L_{0}$ to be a point or $k=0$. The following statements are equivalent.
  1. (a)
    The nearby Lagrangian conjecture holds in $T^{*}L$.
  2. (b)
    Suppose that $L^{\prime}$ is a Lagrangian diffeomorphic to $L$ in a symplectic manifold $M$ and that $L^{\prime}$ is in the Hausdorff closure of $\mathcal{L}{\mathrm{Ham}}(L^{\prime\prime})$ of a $H$-rational Lagrangian $L^{\prime\prime}$ in $M$. Then, $L^{\prime}\in\mathcal{L}{\mathrm{Ham}}(L^{\prime\prime})$. The same holds if $\mathcal{L}{\mathrm{Ham}}(L^{\prime\prime})$ is replaced by $\mathcal{L}{\mathrm{Symp}_{0}}(L^{\prime\prime})$.
@@
Suppose that we are in Case (a). The case of $\mathcal{L}{\mathrm{Symp}_{0}}(L^{\prime\prime})$ follows directly from Proposition 42 together with Theorems 4 and 7. For the case of $\mathcal{L}{\mathrm{Ham}}(L^{\prime\prime})$, take a sequence $\{L_{i}\}$ in that space with limit $L^{\prime}$ diffeomorphic to $L$. By Theorem 7, $L^{\prime\prime}$ is a $H$-rational Lagrangian with same rationality constant as the $L_{i}$’s — the $L_{i}$’s respect the hypotheses of Theorem 4, so that they are $H$-exact in $\mathcal{W}(L^{\prime})$ for $i$ large. Since all $L_{i}$ are Hamiltonian isotopic to each other, their associated symplectic isotopy from Proposition 42 must all have the same flux by Corollary 43. But by that proposition, that flux must tend to 0 as $L_{i}\to L^{\prime}$. Therefore, for $i$ large, there is a symplectic isotopy in $T^{*}L^{\prime}$ sending $L_{i}$ to $L^{\prime}$ with zero flux; again, we suppose that the NLC holds here. By Proposition 2.3 of [Ono08] or Lemma 6.7 of [Sol13], that isotopy must be Hamiltonian, and we have closure, i.e. (a) implies (b).
Suppose that we are in Case (b), and let $L^{\prime\prime}$ be an exact Lagrangian of $T^{*}L$. Then, $L_{i}:=\frac{1}{i}L^{\prime\prime}$ defines a sequence of Lagrangians whose Hausdorff limit is the zero-section $L$. But note that $L_{i}$ is Hamiltonian isotopic to $L^{\prime\prime}$ since it is the image of the exact Lagrangian $L^{\prime\prime}$ by the Liouville flow of $T^{*}L$. Since $\mathcal{L}{\mathrm{Ham}}(L^{\prime\prime})$ is Hausdorff closed, $L\in\mathcal{L}{\mathrm{Ham}}(L^{\prime\prime})$, i.e. the NLC holds in $T^{*}L$.
The Lagrangian $C^{0}$ flux conjecture then follows directly.
Proof of Corollary 9.
Case (b) of Corollary 44 implies the Hausdorff closure of $\mathcal{L}{\mathrm{Ham}}(L^{\prime})$ and $\mathcal{L}{\mathrm{Symp}}(L^{\prime})$ in $\mathcal{L}(L)$ whenever $L^{\prime}$ is $H$-rational and diffeomorphic to $L$. ∎
Remark 12.
Suppose that all exact Lagrangians of $T^{*}L$ are known to be diffeomorphic to the zero-section. Then, the equivalence of (a) and (b) in Corollary 44 actually proves that the NLC for $L$ and the Lagrangian $C^{0}$ flux conjecture for $H$-rational Lagrangians diffeomorphic to $L$ are equivalent. This hypothesis is satisfied when simple homotopy type is enough to determine diffeomorphism type, e.g. when $\dim L\leq 3$, $L=S^{2k+1}/\mathds{Z}_{m}$ for $m\geq 3$ (see [Mil66]), or $L=S^{n}$ for $n\in\{1,2,3,5,6,12,56,61\}$ (see [WX17]).
Remark 13.
@@
5.4 Proof of Proposition 15
We now give a proof of Proposition 15, which appeared in the introduction. The proof is essentially an amalgam of the one of Lemma 13 and of Theorem 5.
Proof of Proposition 15.
As in Lemma 13, take a Riemannian metric $g$ on $M$ which corresponds to a Sasaki metric on $T^{*}L$ on a Weinstein neighbourhood $\mathcal{W}_{r}(L)$ contained in $K$. Let $r_{\mathrm{inj}}(TM|_{K})$ be the injectivity radius of the Riemannian exponential of $g$ restricted to $TM|_{K}$, and take $\delta=\min\{r,r_{\mathrm{inj}}(TM|_{K})\}$. For $\varphi\in\operatorname{Ham}(M)$, we then get a homotopy $\{f_{t}\}$ from $f_{0}=\mathds{1}$ to $f_{1}=\varphi$ using geodesics as in that lemma. Furthermore, by geodesic convexity of $\mathcal{W}_{r}(L)$, $f_{t}(L)\subseteq\mathcal{W}_{r}(L)$ for all $t$ and $f_{1}(L)=\varphi(L)$ is $H$-exact in the neighborhood.
Pick a Hamiltonian isotopy $\{\varphi_{t}\}$ with $\varphi_{1}=\varphi$ and denote by $c:S^{1}\to C^{\infty}(M,M)$ the loop of smooth maps given by the concatenation of $\{f_{t}\}$ with $\{\varphi_{1-t}\}$. Let $x_{0}$ be a fixed point of $\varphi$ such that $t\mapsto\varphi_{t}(x_{0})$ is contractible — this always exists by Floer theory. Then, $f_{t}(x_{0})=x_{0}$ for all $t$ so that $t\mapsto c(t)(x_{0})$ is a contractible loop. But if $x$ is any point in $M$, there is a path $\alpha$ from $x$ to $x_{0}$, so that $(t,s)\mapsto c(t)(\alpha(s))$ defines a free homotopy from the loop defined by $x$ to the one defined by $x_{0}$. In particular, all loops $t\mapsto c(t)(x)$ are contractible.
Given a loop $\gamma$ of $\varphi(L)$, we consider the map $g:\mathds{T}^{2}\to M$ defined by $g(t,s)=c(t)(\gamma(s))$. This decomposes into two cylinders: $\{f_{t}(\gamma(s))\}$ and $\{\varphi_{1-t}(\gamma(s))\}$. Since $f_{t}(L)\subseteq\mathcal{W}_{r}(L)$ for all $t$ and $L$ is exact in $\mathcal{W}_{r}(L)$, the area of the first cylinder is $\lambda_{0}(\gamma)$. But the area of the second cylinder is the flux of the isotopy $\{\varphi_{1-t}\}$ on $\gamma$, which is zero since the isotopy is Hamiltonian. Therefore, $g$ has area $\lambda_{0}(\gamma)$.
If we are in Case (a), then note that the loop $t\mapsto g(t,0)$ is contractible by the above discussion, so that the area of $g$ is actually the area of a sphere. We then conclude similarly to Theorem 5: pick a base $\{\gamma_{i}\}$ of $H_{1}(\varphi(L))^{\mathrm{free}}$, then the corresponding $g_{i}$ respect
$\omega(\operatorname{Ker}c_{1}|_{\pi_{2}(M)})\ni|\omega(g_{i})|=|\lambda_{0}(\gamma_{i})|\leq r\ell_{g}^{\min}(\pi_{*}\gamma_{i})$
by Proposition 41. It thus suffices to take $r$ — and thus $\delta$ — small enough so that $r\max_{i}\ell_{g}^{\min}(\pi_{*}\gamma_{i})$ is smaller than the positive generator of $\omega(\operatorname{Ker}c_{1}|_{\pi_{2}(M)})$. This ensures that $\lambda_{0}|_{\varphi(L)}$ vanishes, i.e. $\varphi(L)$ is exact in $\mathcal{W}_{r}(L)$.
If we are in Case (b), then note that for every $\gamma$, there is some $k$ such that $\gamma^{k}$ is contractible in $M$. In particular, there is a homotopy $\alpha$ from $\gamma^{k}$ to $x_{0}$. Then, $(t,s)\mapsto c(t)(\alpha(s))$ defines a homotopy from $g^{k}$ defined to the loop $t\mapsto\varphi_{t}(x_{0})$, which is contractible by hypothesis. Therefore,
$\lambda_{0}(\gamma)=\frac{1}{k}\lambda_{0}(\gamma^{k})=\frac{1}{k}\omega(g^{k})=0$
for all loops $\gamma$, i.e. $\varphi(L)$ is again exact in $\mathcal{W}_{r}(L)$. ∎
Appendix A Closed embedded loops satisfy Speculation A
Proposition 45
Speculation A holds for any closed embedded curve $L$ in an orientable surface $M$.
@@ Appendix B The case of Lagrangian 2-tori
B.1  Displaceable Lagrangian 2-tori satisfy Speculation B
@@
In particular, Speculation B holds for displaceable rational Lagrangian tori for slightly more general $U$ or in exact symplectic 4-manifolds.
B.2  Non-displaceable Lagrangian 2-tori satisfy Speculation A
@@
Then, Speculation A holds for $L$.
Remark 16.
@@
Appendix C Lagrangian Klein bottles in cotangent bundles
We now focus our efforts on the case of Conjecture C where $L$ is the Klein bottle $K$. This case is already covered by Theorem 2, but we give a more direct, stronger proof of it, which is of independent interest. The proof relies on the deep fact that there is no Lagrangian Klein bottle in $\mathds{C}^{2}$ [She09, Nem09].
Theorem 49
Proof.
Using the flat metric on $\mathds{T}^{2}$ and $K$, the 2:1 covering $p:\mathds{T}^{2}\to K$ lifts to another 2:1 covering $\widetilde{p}:T^{*}\mathds{T}^{2}\to T^{*}K$ which is also a local isometry and symplectomorphism. Therefore, $\widetilde{L}:=\widetilde{p}^{-1}(L)$ must be a (possibly disconnected) Lagrangian submanifold of $D^{*}_{r}\mathds{T}^{2}$. Since $\widetilde{p}|_{\widetilde{L}}$ is also a 2:1 covering, $\widetilde{L}$ must either be two disconnected copies of a Klein bottle or a 2-torus. However, if the former was the case, then each connected component of $\Psi(\widetilde{L})$ would be a Lagrangian Klein bottle in $\mathds{C}^{2}$, which does not exist [She09, Nem09]. Therefore, $\widetilde{L}$ must be a 2-torus. In other words, the composition
[…]
admits a lift to $T^{*}\mathds{T}^{2}$, but the composition
[…]
does not.